OpenStudy (anonymous):

xSquared + 16x + 64 =f(x)

OpenStudy (anonymous):

can you show the work

OpenStudy (anonymous):

does it... it's not set to zero.

OpenStudy (anonymous):

f(x) = (x+8)^2

OpenStudy (anonymous):

x^2+16x+64=f(x) factor by knowing 8+8=16 and 8*8=64 (x+8)^2 x=-8

OpenStudy (anonymous):

-8 will be your zero's

OpenStudy (anonymous):

otherwise, you just want the factor, that is the simplest form of the equation. what does the question ask for?

OpenStudy (anonymous):

it says factor and then solve to find the roots of the function

OpenStudy (anonymous):

ya, so the roots are the zero's and you have the factors

OpenStudy (anonymous):

what would be the final answer

OpenStudy (anonymous):

and the work

OpenStudy (anonymous):

factored form is f(x)=(x+8)^2 roots=-8

OpenStudy (anonymous):

the work is my answer above

OpenStudy (anonymous):

if you want to show every detail in the work, you have to use the quadratic equation, but this one is pretty easy so it isnt entirely necessary

OpenStudy (anonymous):

unless you have just been introduced to the quad eqn, in which case you would use it

OpenStudy (anonymous):

just for good measure, so the prof knows you know what you are doing

OpenStudy (anonymous):

lol i have no idea what im doing tho :/

OpenStudy (anonymous):

then use the quadratic equation and it will give you the roots